OpenStudy (anonymous):

solve the simultaneous equation using the substitution method: x/2+y/4=5 4x/3+5y/8=13

OpenStudy (anonymous):

Q8 of the image

OpenStudy (texaschic101):

multiply the first equation by 4 and the second equation by 24 and get rid of the fractions

OpenStudy (jack1):

x/2 + y/4 = 5 EQN 1 multiply equation 1 by 4 = 2x +y = 20 = EQN 1a re arrange EQN 1a y = 20 - 2x 4x/3 + 5y/8 = 13 EQN 2 sub y = 20 - 2x into EQN 2 4x/3 + 5(y)/8 = 13 4x/3 + 5(20 - 2x)/8 = 13 solve for x

OpenStudy (jack1):

use that x value to solve for y

OpenStudy (jack1):

@harry.g1234 ... how're u going following the above dude...? u ok, want something explained a little clearer...?

OpenStudy (anonymous):

Thanks alot i really appreciate your help its very clear now thanks

OpenStudy (anonymous):

@Jack1

OpenStudy (jack1):

sweet as 'slaters

OpenStudy (anonymous):